OpenStudy (gokuporter):

6-4x=4(3-x)-6 A.3 B.all real numbers C.5 D.8

OpenStudy (alexandervonhumboldt2):

first distribute: 6-4x=4*(3-x)-6 6-4x=12-4x-6 6-4x=6-4x 0=0 thus all real numbers

OpenStudy (gokuporter):

yeah i put b

OpenStudy (alexandervonhumboldt2):

correct